OpenStudy (anonymous):

multiply and simplify: t(t-2)(t+1) Please explain how you did it! Thank you!

OpenStudy (anonymous):

you can first multiply \[(t-2)(t+1)\] together and then multiply the result by t

OpenStudy (anonymous):

\[(t-2)(t+1)=t^2+t-2t-2=t^2-t-2\]

OpenStudy (anonymous):

then \[t(t^2-t-2)=t^3-t^2-2t\]